OpenStudy (anonymous):

A grid shows the positions of a subway stop and your house. The subway stop is located at (–6, 1), and your house is located at (–2, –8). What is the distance, to the nearest unit, between your house and the subway stop?

OpenStudy (johnweldon1993):

You could always use the distance formula \[d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] you have your 2 points (-6,1) (-2,-8) (x1,y1)(x2,y2) plug those in and you'll find your distance

OpenStudy (anonymous):

@FutureMathProfessor Can you help me?

OpenStudy (jim766):

\[\sqrt{(-6--2)^2+(1--8)^2}\] now simplify each parenthese then square each then add them together then take the square root

OpenStudy (jim766):

what is -6 - (-2) ?

OpenStudy (anonymous):

Use the pythagorean theorem. 4^2 + 9^2 = what^2

OpenStudy (anonymous):

@Jim766 It's -4

OpenStudy (anonymous):

-4 times -4 = 16

OpenStudy (jim766):

what is 1 - (-8) =

OpenStudy (anonymous):

9

OpenStudy (anonymous):

9*9=81

OpenStudy (jim766):

so now we have \[\sqrt{16 + 81}\] add them together

OpenStudy (anonymous):

Ohh, okay! 97 squared is 9.8

OpenStudy (jim766):

\[\sqrt{97} \approx 9.85\]

OpenStudy (jim766):

right,

OpenStudy (jim766):

you just plug it in and then do what it says

OpenStudy (anonymous):

Awesome, thank you.

OpenStudy (jim766):

yw